Joey makes 100 pounds of a coffee mix. Columbian coffee selling for $8.75 per pound is mixed with Brazilian coffee selling for $3.75 per pound. The final mix sells for $6.35 per pound. How much of each type of coffee is used in the mix?
Answer by mananth(16949)

cost-------------------- quantity
8.75 ---------------- x
3.75 ---------------- 100-x
6.35 ---------------- 100
...
8.75x+3.75(100-x)=6.35* 100
8.75x+375-3.75x= 635
8.75x-3.75x=635-375
5x=260
/5
x=52 pounds of $8.75
48 pounds of $3.75%